OpenStudy (ziporah):

8x - 11 = -2x + 15 -5. What steps would you take to solve this?

OpenStudy (3mar):

\[8x-11=-2x+15-5\] \[8x-11=-2x+10\] \[8x+2x=10+11\] \[10x=21\] \[x=2.1\] Hope that helps